Circular No. 5836

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J1008–57 M. T. Stollberg, M. H. Finger, R. B. Wilson, B. A. Harmon, B. C.

Rubin, N. S. Zhang, and G. J. Fishman report for the Compton Gamma Ray Observatory BATSE team: “Hard x-ray pulsations in the energy range 20-120 keV are being detected from a previously uncatalogued source in the Carina-Vela region. The source location, determined from the observed times of Earth occultation, is centered at α = 10^{h}08^{m}, δ = −57^{o}.5 (equinox 2000.0). The estimated 1σ errors are bounded by a box with the following corners: α = 10^{h}09^{m}.5, δ = −57^{o}24′; 10^{h}03^{m}.5, −56^{o}24′; 10^{h}07^{m}.0, −57^{o}30′; 10^{h}12^{m}.5, −58^{o}42′. The pulse period on July 19 was 93.587 ± 0.005 s, corrected to the solar system barycenter. The present outburst was first detected on July 14. The pulsed flux in the range 20-50 keV increased from a detection threshold of ∼ 50 mCrab to ∼ 1.1 Crab (pulsed) on July 18, and this flux has continued to increase slightly since that time. The spectrum of the phase-averaged pulsed flux between E = 20 and 100 keV on July 23-24 has been fitted by an optically thin thermal bremsstrahlung model of the form A/E exp(−E/kT), with A = 0.21 ± 0.02 photon cm^{−2} s^{−1} and kT = 18.9 ± 0.9 keV.”

NOVA OPHIUCHI 1993 C. E. Woodward, University of Wyoming; M. A. Greenhouse, National

Air and Space Museum; and D. Van Buren, California Institute of Technology, write: “Spectroscopy of N Oph 1993 obtained on June 2.3 UT (62 days after outburst; cf. IAUC 5765) with the Kitt Peak National Observatory 2.1-m telescope (+ CRSP) reveals the following infrared coronal lines (wavelengths listed are observed values, with uncertainty in the last digit given in parentheses): [Ca VIII] 2.322(2)-μm (2_{1/2}^{P}-2_{3/2}^{P}) and [Si VII] 2.470(2)- μm (3_{2}^{P} -3_{1}^{P}). Emission from hydrogen (Brγ 2.169(1)-μm) and helium (He I 2.060(2)-μm) also were present in the spectra. The observed integrated intensities were (in units of 10^{−19} W/cm^{2}): Brγ, 1.06 ± 0.01; [Ca VIII], 0.26 ± 0.05; [Si VII], 0.65 ± 0.13. The lines were velocity-resolved at ∼ 3600 km/s FWHM. The rapid onset of coronal line emission in N Oph 1993 is similar to that observed in V1500 Cyg and V446 Her. Further spectroscopy is encouraged.”

1993 July 27 Daniel W. E. Green